OpenStudy (anonymous):

What is the slope of the line that passes through (3, –7) and (–1, 1)?

OpenStudy (dumbcow):

\[m = \frac{y2 - y1}{x2-x1} = \frac{-7-1}{3-(-1)} = \frac{-8}{4} = -2\]